OpenStudy (anonymous):

HELP solve for x? 3/(x-3)=x/(x-3)-3/2 and is it extraneous or non extraneous?

OpenStudy (anonymous):

what are you trying to solve ???

OpenStudy (anonymous):

@TrapBoiTrynaEducaate I edited it

OpenStudy (anonymous):

3/(x-3) - x/(x-3) = - 3/2 (3 - x) / (x - 3) = - 3/2 3 - x = -3/2 (x-3) 3 - x = -(3/2)x + 9/2 6 - 2x = -3x + 9 x = 3

OpenStudy (anonymous):

@TrapBoiTrynaEducaate is the solution extraneous or non extraneous?

OpenStudy (anonymous):

idek i just did it like that

OpenStudy (cwrw238):

try plugging in x = 3 into original equation

OpenStudy (anonymous):

@TrapBoiTrynaEducaate if you plug in x=3 you get 0 in the denominator, so I guess it's extraneous?

OpenStudy (anonymous):

yup

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

@cwrw238 thank you

OpenStudy (cwrw238):

yw